# Training LLMs to Better Self-Debug and Explain Code

### v1

@@ 4.4 RQ4: Quality of generated explanation
Raters Prompting SFT RL Prompting SFT RL Prompting SFT RL Prompting
@@
Due to the effort required by human rating, we only let developers rate the StarCoder models’ and GPT-3.5-Turbo’s generations and each explanation is rated as the average of scores given by two developers. The human rating is consistent with GPT-4 indicating that SFT improves the explanation quality over prompting and RL training takes code explanation semantics into consideration of rewards, thus bringing continuous improvement to the explanation generation. Rubrics and examples of human rating are shown in Appendix A.6.
@@ 5 Limitation
One potential limitation of our study is the reliance on specific large language models (LLMs) from which we collect code explanation and refinement data. Our automated framework is designed to be independent of any specific LLM, and for this study, we use GPT-3.5 and CodeLlama-34B to collect training data, and both bring significant improvement through SFT and RL. However, for future work, it would be interesting to explore the use of other LLMs, including smaller models or a mix of diverse LLMs, to gather explanation and refinement data.
@@ 6.1 Large Language Models for Code
6.2 Self-Debugging
Existing techniques have studied the possibility of using LLMs to refine their generations. Yet, most techniques are prompting LLMs with execution results [13, 14, 42, 43, 44, 45, 46, 47] for the refinement. Such prompting approaches bring limited improvement to smaller open-sourced LLMs compared to GPT-3.5. Other techniques train LLMs to self-debug. ILF [48] uses human-annotated feedback information and thus is unscalable, CYCLE [49] and Self-Edit [50] use SFT to fine-tune LLM to generate the refinement only based on the unit test execution feedback.
This work has three differences compared with others that train LLMs: (1) we train LLMs to generate code explanation followed by refinement, which provides additional information to users, (2) we do not require human-annotated training data but propose a scalable pipeline to automatically collect and verify data from another LLM, (3) we design novel reward functions in the RL training stage, considering both the code and explanation quality, which beings extra improvement.
@@ 7 Conclusion
This work shows the necessity of training open-sourced LLMs to self-debug and proposes a scalable framework consisting of automated data collection, data verification, supervised fine-tuning, and reinforcement learning with new reward designs, to improve LLMs’ self-debugging ability. The data collection is independent of LLMs. Our experiments using either GPT-3.5-Turbo or CodeLlama-34B achieve great improvements. The data verification ensures the high quality of the collected code explanation and refinement data. SFT on our collected data brings a significant boost to LLMs’ self-debugging ability, resulting in up to 15.92% higher pass@1, 9.30% higher pass@10, and 10.15% higher success refinement rate. RL training using our novel reward design further brings up to 3.54% higher pass@1, 2.55% higher pass@10, and 3.62% higher success refinement rate. Comprehensive analytical experiments validate the generalizability of our approach and illustrate the iterative refinement ability of the trained LLMs. The human evaluation indicates that the explanations generated by the LLMs trained with our approach are of higher quality and can effectively serve developers for debugging purposes.
@@ A.1 Limited refinement ability of open-sourced LLMs
We test two prompts of querying LLMs for code refinement, letting LLMs generate the code refinement only (Figure 5 (3.1)), or letting LLMs first explain why the code is wrong and then generate the refinement (Figure 5 (3.2)).
Table 8 shows the success rate of refinement generated by StarCoder-15B, CodeLlama-7B, CodeLlama-13B, and GPT-3.5 using greedy decoding. Mostly, when LLMs explain the wrong code first, they are more likely to provide the correct refinement. Yet, all the three open-source LLMs we test, even with the best efforts, show limited code refinement ability. For instance, StarCoder-15B is only able to refine 4.43 – 5.10% of the wrong code, which motivates the training of open-sourced LLMs to obtain code explanation and refinement ability.
@@
StarCoder-15B 2.58% 4.43% 2.04% 5.10%
CodeLlama-7B 7.42% 6.71% 5.32% 8.52%
CodeLlama-13B 9.80% 10.20% 5.88% 8.24%
GPT-3.5 26.24% 28.90% 32.32% 33.33%
@@ A.2 Data collection and data format for training
Figure 6 shows an example of the prompt we used to collect code explanation and refinement training data from GPT-3.5 (same for collecting data from CodeLlama-34B). We use a temperature of 0.8 to let GPT-3.5 sample 10 code explanations and refinements per prompt.
@@
The data collected contains both the code explanation and refinement, and we construct two formats of instruction-tuning data for SFT. Figure 7 (left) shows the instruction asking LLMs for code refinement only, where the instruction contains the task description, wrong initial solution, and execution feedback. LLMs are trained to generate the output refinement given the instruction. Figure 7 (right) shows another instruction asking LLMs for both code explanation and refinement. We include both types of instruction data in the training, which increases the training data diversity and enhances LLMs’ robustness against different prompts.
@@ A.4.1 Comparison with SFT on code generation
Table 9 shows the comparison between SFT CodeLlama-7B trained with our collected data (including the code generation, code explanation, and code refinement data), and the CodeLlama-7B trained with code generation data only. We find that although training with code generation enables LLMs to get comparable (or even higher pass@k on HumanEval) pass@k on one round of code generation, the LLMs cannot obtain strong self-debugging ability from code generation data. The LLM trained with code generation data can only improve the pass@k very little after self-debugging. By contrast, the LLM trained with the full collected data gets significantly higher pass@k after refinement, showing its strong self-debugging ability.
@@
SFT on code Refine 49.18 72.78 42.91 74.57 43.13 61.81 36.12 65.77
generation data Expl. + Refine 49.58 73.07 43.88 75.19 43.58 61.69 36.37 66.24
@@
SFT on the Refine 58.07 77.34 52.65 80.71 51.64 71.04 46.61 74.43
full collected data Expl. + Refine 57.98 77.92 52.98 82.22 51.55 70.94 47.62 75.54
@@ A.4.3 Generalizability
Table 11: Overall pass@k on MBPP & HumanEval and MBPP+ & HumanEval+, trained with CodeLlama-34B’s data.
CodeLlama MBPP & HumanEval MBPP+ & HumanEval+
13B SFT RL SFT RL
Refine Expl. + Ref. Refine Expl. + Ref. Refine Expl. + Ref. Refine Expl. + Ref.
pass@1 57.85 (+9.13) 58.10 (+9.14) 58.30 (+0.45) 58.71 (+0.61) 50.46 (+7.74) 51.43 (+8.68) 57.26 (+6.80) 56.45 (+5.02)
pass@10 80.61 (+3.47) 80.84 (+3.27) 81.47 (+0.86) 81.47 (+0.63) 75.09 (+3.92) 75.5 (+3.65) 77.37 (+2.28) 76.92 (+1.42)
Table 12: Success refinement rate of different approaches over four benchmarks, trained with CodeLlama-34B’s data.
@@
Prompt. SFT RL Prompt. SFT RL
CodeLlama-13B 11.64 15.60 (+3.96) 22.74 (+7.14) 11.97 16.83 (+4.86) 21.61 (+4.78)
Table 12 shows the success refinement rate generated by CodeLlama-13B trained with data collected from CodeLlama-34B. The result is averaged over four benchmarks. SFT successfully refines 15.60 – 16.83% wrong solutions, outperforming the prompting approach by 3.96 – 4.86%. RL training further increases the refinement rate by 4.78 – 7.14% over SFT. We note that using CodeLlama-34B’s data, the improvement brought by SFT is smaller compared with training with GPT-3.5-Turbo’s data, which may be due to the slightly lower quality of code explanation and refinement data from CodeLlama-34B. Yet, RL training brings the refinement rate to a similar level to that trained with GPT-3.5-Turbo’s data, suggesting that RL training can potentially overcome the quality difference from open-sourced LLMs to commercial LLMs.
@@ A.7 Potential impact
Using LLMs to help coding is popular nowadays. This work proposes a technique to train LLMs to explain and self-refine code, which aims to improve developers’ coding experience with LLMs. We also call for training LLMs to take feedback beyond the prompting approach to improve the LLMs’ self-debugging ability.
The technique is supposed to be on the same track as all the existing LLMs for code generation. Thus, we think no special concerns about broader impact need to be highlighted here.

### v2

@@ 4.3.1 Data collection using open source LLM
Raters Prompt. SFT RL Prompt. SFT RL Prompt. SFT RL Prompt.
@@
4.3.2 Data Collection Using Self-Bootstrap
We also investigate the feasibility of using an LLM to self-bootstrap its training data, specifically by using CodeLlama-7B to generate the data that is then used for its own SFT and RL training.
Table 7 presents the pass@k results of CodeLlama-7B trained with self-bootstrapped data, with Table 8 showing the overall comparison. Compared to prompting, SFT with self-bootstrapped data still delivers up to 8.25% and 2.14% improvements in pass@1 and pass@10 on MBPP and HumanEval, and up to 5.33% and 0.55% improvements on pass@1 and pass@10 on MBPP+ and HumanEval+. Additionally, RL training using the self-bootstrapped data results in a further 0.71% improvement on MBPP and HumanEval, and up to a 0.78% increase on MBPP+ and HumanEval+. These findings suggest that while self-bootstrapped data enables SFT to provide substantial gains over prompting, RL training offers less improvement compared to using data from stronger LLMs, such as CodeLlama-34B or GPT-3.5-Turbo.
4.4 RQ4: Quality of generated explanation
We assess whether explanations for incorrect code are useful for developers in understanding their bugs. To do this, we randomly sample 50 problems with initial incorrect solutions from the MBPP and HumanEval benchmarks and use different LLMs to generate explanations for the wrong code. Each explanation is scored on a scale from 1 to 5, based on its correctness and helpfulness, where 1 indicates a completely incorrect or misleading explanation, and 5 denotes a correct explanation that also provides a detailed hint on how to fix the code. Both GPT-4 and human developers are used as evaluators. For GPT-4, we follow prior work [27] and prompt it to score each explanation. The results are presented in Table 9. Both SFT and RL lead to improved explanation quality compared to prompting, with GPT-4 assigning higher scores to models trained using our approach. Notably, the gap between GPT-3.5-Turbo and the trained LLMs significantly narrows after fine-tuning.
Given the time required for human evaluation, we only asked developers to rate explanations from StarCoder models and GPT-3.5-Turbo. Each explanation is scored by two developers, and their ratings are averaged. The human evaluations align with GPT-4’s, confirming that SFT improves explanation quality over prompting, while RL further enhances explanations by incorporating code explanation semantics into the reward design. Detailed rubrics and examples of human evaluations can be found in Appendix A.6.
@@ 5 Limitation
One potential limitation of our study is the reliance on specific large language models (LLMs) from which we collect code explanation and refinement data. Our automated framework is designed to be independent of any specific LLM, and for this study, we use GPT-3.5, CodeLlama-34B, and CodeLlama-7B itself to collect training data, and both bring significant improvement through SFT and RL. However, for future work, it would be interesting to explore the use of other LLMs, including smaller models or a mix of diverse LLMs, to gather explanation and refinement data.
@@ 6.1 Large language models for code
6.2 Self-debugging and self-refinement
Existing techniques have studied the possibility of using LLMs to refine their generations. Yet, most techniques are prompting LLMs with execution results [13, 14, 42, 43, 44, 45, 46, 47, 48] for the refinement. Such prompting approaches bring limited improvement to smaller open-sourced LLMs compared to GPT-3.5. Other techniques train LLMs to self-debug. ILF [49] uses human-annotated feedback information and thus is unscalable, CYCLE [50] and Self-Edit [51] use SFT to fine-tune LLM to generate the refinement only based on the unit test execution feedback. OpenCodeInterpreter [52] and EURUS [53] construct high-quality multi-turn interaction datasets using GPT-3.5-Turbo and GPT-4 to fine-tune LLM for self-refinement.
This work has four differences compared with others that train LLMs: (1) we train LLMs to generate code explanation followed by refinement, which provides additional information to users, (2) we do not require human-annotated training data but propose a scalable pipeline to automatically collect and verify data from another LLM, (3) our data collection pipeline can be generalized to open-sourced LLM or even the same LLM itself, and (4) we design novel reward functions in the RL training stage, considering both the code and explanation quality, which beings extra improvement.
@@ 7 Conclusion
This work highlights the importance of training open-source LLMs to self-debug and introduces a scalable framework that includes automated data collection, verification, supervised fine-tuning, and reinforcement learning with novel reward designs to enhance LLMs’ self-debugging capabilities. Our data collection process is model-agnostic, as demonstrated by the improvements achieved with both GPT-3.5-Turbo and CodeLlama. The data verification ensures high quality of code explanations and refinements. Fine-tuning on this data significantly boosts the LLMs’ self-debugging abilities, yielding up to a 15.92% increase in pass@1, a 9.30% increase in pass@10, and a 10.15% increase in successful refinements. Reinforcement learning, utilizing our novel reward design, further enhances performance, with additional gains of up to 3.54% in pass@1, 2.55% in pass@10, and 3.62% in successful refinement rates. Comprehensive analytical experiments confirm the generalizability of our approach and demonstrate the iterative refinement capabilities of the trained models. Moreover, human evaluations indicate that the LLMs trained with our framework produce higher-quality explanations, effectively aiding developers in understanding and resolving bugs in source code.
@@ A.1 Limited refinement ability of open-sourced LLMs
We evaluate two prompting strategies for querying LLMs to perform code refinement: one where the LLMs directly generate the refinement (Figure 5 (3.1)), and another where they first explain why the code is incorrect before generating the refinement (Figure 5 (3.2)). Table 10 reports the success rates of refinements generated by StarCoder-15B, CodeLlama-7B, CodeLlama-13B, and GPT-3.5 using greedy decoding. Generally, LLMs that first explain the incorrect code are more likely to generate accurate refinements. However, despite these efforts, all three open-source LLMs exhibit limited refinement capabilities. For example, StarCoder-15B successfully refines only 4.43% to 5.10% of incorrect code, underscoring the need for training open-source LLMs to enhance their ability to explain and refine code.
@@
StarCoder-15B 2.58 4.43 2.04 5.10
CodeLlama-7B 7.42 6.71 5.32 8.52
CodeLlama-13B 9.80 10.20 5.88 8.24
GPT-3.5 26.24 28.90 32.32 33.33
@@ A.2 Data collection and data format for training
Figure 6 provides an example of the prompt we used to collect code explanation and refinement training data from GPT-3.5 (the same approach was used for CodeLlama-34B). We set the temperature to 0.8, allowing GPT-3.5 to generate 10 code explanations and refinements per prompt. The collected data includes both the explanation of the incorrect code and the corresponding refinement. From this, we construct two formats of instruction-tuning data for supervised fine-tuning (SFT).
Figure 7 (left) shows an instruction that asks LLMs to provide only the code refinement, where the instruction includes the task description, the incorrect initial solution, and execution feedback. The LLMs are trained to generate the refined solution based on this input. Figure 7 (right) presents a different instruction that asks for both the code explanation and the refinement. By incorporating both types of instruction data during training, we increase the diversity of the training set and improve the LLMs’ robustness across different prompts.
@@ A.4.1 Comparison with SFT on code generation
Table 11 shows the comparison between CodeLlama-7B trained with LeDex, and the CodeLlama-7B trained with code generation data only. Although training with code generation enables LLMs to get comparable (or even higher pass@k on HumanEval) pass@k on one round of code generation, the LLMs cannot obtain strong self-debugging ability from code generation data. The LLM trained with code generation data can only improve the pass@k very little after self-debugging. By contrast, the LLM trained with the full collected data (code generation, code explanation, and code refinement data) gets significantly higher pass@k after refinement, showing its strong self-debugging ability.
@@
Only Code Generation Data Refine 49.18 72.78 42.91 74.57 43.13 61.81 36.12 65.77
Expl. + Refine 49.58 73.07 43.88 75.19 43.58 61.69 36.37 66.24
@@
LeDex’s Full Data Refine 58.07 77.34 52.65 80.71 51.64 71.04 46.61 74.43
Expl. + Refine 57.98 77.92 52.98 82.22 51.55 70.94 47.62 75.54
@@ A.4.3 Generalizability
Table 13: Overall pass@k on MBPP & HumanEval and MBPP+ & HumanEval+, trained with CodeLlama-34B’s data. Blue numbers show the improvement.
CodeLlama-13B MBPP & HumanEval MBPP+ & HumanEval+
LeDex SFT LeDex RL LeDex SFT LeDex RL
Refine Expl. + Refine Refine Expl. + Refine Refine Expl. + Refine Refine Expl. + Refine
pass@1 57.85 +9.13 58.10 +9.14 58.30 +0.45 58.71 +0.61 50.46 +7.74 51.43 +8.68 57.26 +6.80 56.45 +5.02
pass@10 80.61 +3.47 80.84 +3.27 81.47 +0.86 81.47 +0.63 75.09 +3.92 75.5 +3.65 77.37 +2.28 76.92 +1.42
Table 14: Success refinement rate of different approaches over four benchmarks, trained with CodeLlama-34B’s data. Blue numbers show the improvement.
@@
Prompt. LeDex SFT LeDex RL Prompt. LeDex SFT LeDex RL
CodeLlama-13B 11.64 15.60 +3.96 22.74 +7.14 11.97 16.83 +4.86 21.61 +4.78
Table 14 presents the success refinement rates achieved by CodeLlama-13B when trained on data collected from CodeLlama-34B, averaged across four benchmarks. SFT refines 15.60% to 16.83% of incorrect solutions, outperforming the prompting approach by 3.96% to 4.86%. RL training further boosts the refinement rate, improving it by 4.78% to 7.14% over SFT. Notably, the improvement from SFT is somewhat smaller when using CodeLlama-34B’s data compared to GPT-3.5-Turbo’s data, likely due to the slightly lower quality of explanations and refinements generated by CodeLlama-34B. However, RL training raises the refinement rate to levels comparable to those achieved with GPT-3.5-Turbo’s data, indicating that RL training can mitigate the quality differences between data generated by open-source LLMs and commercial LLMs.
@@ A.7 Potential impact
Using LLMs to help coding is popular nowadays. This work proposes a technique to train LLMs to explain and self-refine code, which aims to improve developers’ coding experience with LLMs. We also call for training LLMs to take feedback beyond the prompting approach to improve the LLMs’ self-debugging ability. The technique is supposed to be on the same track as all the existing LLMs for code generation. Thus, we think no special concerns about broader impact need to be highlighted here.